You're working for Ford as a safety engineer. You job is to check the stopping ability of different types of brakes. To do this, you measure the stopping distance of cars in order to determine their acceleration. If a car moving at an initial velocity of 25.6 m/sec has a stopping distance of 99.6 meters, what is the magnitude of it's acceleration?

Give you answer in units of m/s2 to at least three significant figures to avoid being counted off for rounding.
Answer #1

using 3rd equation of motion

v^2 = u^2 - 2 a d

0 = 25.6^2 - 2 * a* 99.6

a = 3.29 m/s^2

==========

do comment in case any doubt, will reply for sure. Goodluck